### v1

# Parameter estimation of microlensed gravitational waves with Conditional Variational Autoencoders

Abstract
Gravitational lensing of gravitational waves (GWs) provides a unique opportunity to study cosmology and astrophysics at multiple scales. Detecting microlensing signatures, in particular, requires efficient parameter estimation methods due to the high computational cost of traditional Bayesian inference. In this paper we explore the use of deep learning, namely Conditional Variational Autoencoders (CVAE), to estimate parameters of microlensed binary black hole (simulated) waveforms. We find that our CVAE model yields accurate parameter estimation and significant computational savings compared to Bayesian methods such as bilby (up to five orders of magnitude faster inferences). Moreover, the incorporation of CVAE-generated priors in bilby reduces the average runtime of the latter in about 48% with no penalty on its accuracy. Our results suggest that a CVAE model is a promising tool for future low-latency searches of lensed signals. Further applications to actual signals and integration with advanced pipelines could help extend the capabilities of GW observatories in detecting microlensing events.
I Introduction
The first three observing runs of the network of gravitational-wave (GW) detectors Advanced LIGO [1], Advanced Virgo [2] and KAGRA [3] have led to the detection of 90 compact binary coalescence (CBC) events [4, 5, 6, 7]. Their analysis is impacting different fields of research, e.g. astronomy, astrophysics, cosmology, and fundamental physics, in a significant way. In particular the wealth of new data has provided updated information on the CBC merger rate and has allowed for new studies of populations of compact objects (specifically regarding the distribution of black hole masses) and for new tests of general relativity [8, 9]. One distinct aspect being scrutinized is the possibility of finding evidence of gravitational lensing on the observed GW signals [10, 11]. If it took place, lensing by massive objects along the line of sight between the detector and the source might significantly distort the signal leading to e.g. beating patterns, signal amplitude changes, signal repetition, or a combination of these phenomena.
The LIGO-Virgo-KAGRA (LVK) Collaboration has conducted several searches for GW lensing, including both strong lensing and microlensing signatures. The most recent searches have accounted for all signals from the whole sample of binary black hole (BBH) mergers reported in the third observing run [10, 11]. None of the searches conducted in those studies has yielded conclusive evidence for GW lensing [12]. Other groups, however, have put forward different proposals suggesting that the high masses observed in some BBH systems could be explained by the magnification effect of gravitational lensing [13, 14, 15] (see also related discussions in [10, 16]).
Despite lensing of gravitational waves has not yet been established, it is firmly regarded as a new instrument for GW astrophysics with a remarkable potential for new discoveries, as highlighted in studies related to the LISA mission [17]. Certain gravitational lenses might be composed of objects that neither emit light nor neutrinos, making them undetectable through conventional means. Observing GW lensing could hence provide crucial insights into populations of elusive or poorly understood objects, such as intermediate-mass black holes [18, 19, 20], dark matter haloes [21, 22, 23], cosmic strings [24, 25, 26], or even potentially reveal entirely new classes of objects. The first detection of gravitational lensing by such phenomena would provide direct evidence of compact dark matter objects, offering a groundbreaking advance in astrophysics [27, 28, 21, 29, 30, 22, 31, 32, 33, 34, 35, 36].
Just as it happens for electromagnetic signals, the effects and impact of lensing on a GW can be remarkably diverse. On the one hand, lensing amplification can cause the source’s distance and mass to be underestimated, leading to incorrect estimation of parameters. If a GW is lensed, it might appear closer and more massive than it truly is, which makes the identification of the effect crucial to avoid such errors. On the other hand, strong lensing, caused by galaxies, would produce repeated GW events separated by minutes to months. This phenomenon occurs when large masses significantly bend spacetime, creating multiple images of the source [37, 38, 39]. In contrast, microlensing, caused by stars or compact objects, results in much shorter time delays, with duplicate events separated by milliseconds to seconds. Unlike strong lensing, microlensing can magnify signals or create a “beating” pattern in the waveform, where the time delay between paths is comparable to the wave’s period.
The investigation of the lensing phenomenon has received significant attention in GW astronomy, with an increasing body of work that offers unique insights into small-scale structures in the universe [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74]. In this paper we report on our contribution to these efforts by discussing the results from a study on the microlensing case, particularly in the point-mass lens model. This is the simplest and most illustrative model for a microlensing object, where the lensing effects depend only on the redshifted lens mass $M_{Lz}$ and on the parameter $y$, which defines the position of the source relative to the line of sight.
The straightforward approach to detecting microlensing is to do Bayesian parameter estimation of individual events with lensing added to the waveform model. Currently, the pipeline in charge of performing these searches is Gravelamps [59]. The biggest limitation of a Bayesian approach is computational efficiency, as it can be extremeley time consuming. Recent studies [75, 76] have demonstrated that deep-learning techniques, particularly Conditional Variational Autoencoders (CVAE) [77], offer a powerful way to perform fast parameter estimation in (unlensed) GW data. Indeed, CVAEs have shown impressive potential in capturing complex features of GW signals, making them a valuable tool in contexts where speed is essential (e.g. low-latency searches). Other deep-learning approaches have been used for the identification of microlensed [78] and strong-lensed [79, 80] signals, as well as on parameter estimation and detection of strong lensing in the electromagnetic case [81, 82]. Despite these advances, there are currently no studies that apply deep-learning techniques specifically to the case of parameter estimation of microlensed gravitational waves. The aim of the present investigation is to bridge that gap, leveraging the unique capabilities of CVAEs to address the complex problem of parameter estimation in microlensing scenarios.
To do so we train a CVAE on a dataset of simulated, microlensed GW signals, enabling it to learn a probabilistic latent representation of the lensing parameters conditioned on the wave’s time series. The results show that such a model could be deployed in inference tasks, offering a practical way to integrate microlensing models into real-time analysis frameworks, complementing traditional Bayesian parameter estimation methods, such as those based on packages like Bilby [83] or Gravelamps [59]. In particular, we demonstrate that incorporating CVAE-generated priors into Bilby significantly reduces its inference time, thereby enhancing the practicality of Bayesian methods for microlensing detection.
This paper is organized as follows: In Section II we review the theoretical framework in which microlensing and, in particular, the point-mass lens model, are based. Section III discusses the methodology employed in our study, this is, CVAEs. Next, in Section IV, we present the different datasets used, together with some specifications on the data generation process. Section V presents the results of the training and testing processes, as well as a time comparison between CVAEs and Bilby. Finally, the conclusions of this work are presented in Section VI. Specific details concerning the architecture of the model, the training and testing processes and the comparison with Bilby are reported in Appendix A.
II Microlensing
Any astrophysical object with sufficient mass can act as a lens for a passing gravitational wave. Lensing effects occur when the source, the lens and the observer are all aligned within the Einstein angle $\theta_{\rm E}=R_{\rm E}/d_{L}$, i.e., the lens is located near the line of sight. The Einstein radius $R_{\rm E}$ is commonly expressed as:
$R_{\rm E}=\sqrt{\frac{4GM_{L}}{c^{2}}\,\frac{d_{LS}\,d_{L}}{d_{S}}},$ (1)
where $M_{L}$ is the mass of the lens, $d_{LS}$ is the angular diameter distance between the source and the lens, and $d_{L}$ and $d_{S}$ are the angular diameter distances to the lens and source at redshifts $z_{L}$ and $z_{S}$, respectively. $R_{\rm E}$ represents the characteristic length scale on the lens plane, is proportional to $\sqrt{M_{L}}$ and is typically much smaller than the cosmological distances $d_{LS}$, $d_{L}$, and $d_{S}$. This enables the lens mass to be projected onto a lens plane. In the thin-lens approximation, gravitational waves propagate freely outside the lens, interacting only with a two-dimensional gravitational potential at the lens plane, where the lensing effect is ultimately captured in the transmission factor $F$ [84].
An unlensed GW signal from the source can be described by its frequency-domain strain $\tilde{h}(f)$, which is the Fourier transform of the time-domain strain $h(t)$. After the signal passes through a lens, the resulting lensed waveform $\tilde{h}_{L}(f)$, which is ultimately detected, is the product of the transmission factor $F(f)$ and the original unlensed waveform:
$\tilde{h}_{L}(f)=F(f)\cdot\tilde{h}(f).$ (2)
The transmission factor is determined by the Fresnel-Kirchhoff diffraction integral across the lens plane [84]
[…]
where the lensing time delay function is given by
$t_{d}(\mathbf{x},\mathbf{y})=t_{M}\,\left(\frac{1}{2}\,|\mathbf{x}-\mathbf{y}|^{2}-\psi(\mathbf{x})+\phi_{m}(\mathbf{y})\right)$ (4)
with the characteristic lensing time in physical units
$t_{M}=(1+z_{L})\,\frac{\xi_{0}^{2}\,d_{S}}{c\,d_{L}d_{LS}}.$ (5)
Here, $\xi_{0}$ is an arbitrary length scale in the lens plane, used to normalize both the impact parameter $\xi$ and the source position $\eta$ as follows [42]: $\mathbf{x}=\mbox{\boldmath{$\xi$}}/\xi_{0}$, $\mathbf{y}=\mbox{\boldmath{$\eta$}}\,d_{L}/(\xi_{0}d_{S})$. In this notation, $\eta$ indicates the position of the source in the source plane, while $\mathbf{y}$ represents the projection of the source onto the lens plane, normalized by $\xi_{0}$. For axially symmetric lenses, the vector $\mathbf{y}$ can always be reduced to the scalar $y$ by properly rotating the coordinate axes. The factor $(1+z_{L})$ is included to account for cosmological distance, where $z_{L}$ represents the redshift of the lens. The properties of the lens are captured by the lensing potential $\psi(\mathbf{x})$. In the absence of lensing, where $\psi(\mathbf{x})=0$, the transmission factor (3) simplifies to $|F|=1$, leaving the waveform unaffected. For convenience, the phase $\phi_{m}(\mathbf{y})$ is defined such that the minimum of $t_{d}(\mathbf{x},\mathbf{y})$ for a fixed $\mathbf{y}$ is zero.
Next, we will consider an isolated point mass as the lens. In gravitational lensing, the point mass lens model (PML) is appropriate when the physical size of the lens is much smaller than the Einstein radius, as in the case of black holes, dense dark matter clumps, and similar compact objects. Due to its simplicity, the PML model has been widely used in the literature to interpret both electromagnetic [85, 86, 84, 87] and GW lensing [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73].
For the PML model, a natural choice for the length scale $\xi_{0}$ is the Einstein radius. With this normalization, the lensing potential is $\psi(\mathbf{x})=\ln{|\mathbf{x}|}$, for which the diffraction integral (3) can be solved analytically [84, 85]
[…]
where $\nu\equiv ft_{M}$ is the frequency of the gravitational wave $f$ scaled by the characteristic time $t_{M}$, $\Gamma(z)$ is the Gamma function, and ${}_{1}F_{1}(a,b,z)$ is the confluent hypergeometric function [^1] In our code we take the complex conjugate by replacing $\mathrm{i}$ with $-\mathrm{i}$ to align with the Fourier transform convention used in the Python libraries of the LVK Collaboration [88].. With the chosen normalization $\xi_{0}=R_{\rm E}$, the time $t_{M}$, as expressed by Eq. (5), is proportional to the mass of the lens:
[…]
where ${\cal R_{\rm S}}=2GM_{Lz}/c^{2}$ denotes the Schwarzschild radius, and $M_{Lz}=M_{L}\,(1+z_{L})$ is the redshifted mass of the lens.
The absolute value of the transmission factor is obtained from Eq. (6) as follows [85]
[…]
As seen from Eqs. (6) and (8) the transmission factor, which is a function of frequency $f$, depends on two parameters: (i) the mass of the lens $M_{Lz}$ through the time $t_{M}$ and (ii) the scaled offset of the source $y$. Its behavior has already been described in the literature [85, 42, 43] (in current notation, see also Refs. [61, 65]).
We are interested in the behavior of the transmission factor within the frequency range detectable by the LVK network, which spans approximately from 30 Hz to 1 kHz [89, 90, 91, 92, 93, 94]. For a given choice of parameters $M_{Lz}$ and $y$, the lensing effect on GWs in the LVK range can fall into the amplification region, the oscillating region, or an intermediate region, depending on the lens mass [61]. For lower masses (e.g. $M_{L}=30M_{\odot}$), the primary effect is amplification. As the mass increases, oscillatory effects with a beating pattern emerge in the waveform [65]. At high frequencies, calculating the transmission factor (8) in the oscillating region becomes numerically expensive. Fortunately, in this limit, the dominant contribution arises from two well-defined images of the source, corresponding to two stationary points of the time delay function. The transmission factor in this geometrical optics (GO) limit can then be written as a sum over these stationary points [84, 41, 42],
$F_{\rm GO}=\sum_{j}|\mu_{j}|^{1/2}e^{\mathrm{i}(2\pi f\,t_{d}(\mathbf{x}_{j},\mathbf{y})-n_{j}\pi/2)},$ (9)
where $\mu_{j}=(\det\left(\partial\mathbf{y}/\partial\mathbf{x}_{j}\right))^{-1}$ is the magnification of the j-th image and $n_{j}=0,1,2$ corresponds to the Morse index for a minimum, saddle point and maximum of $t_{d}$, respectively. The positions of the images $\mathbf{x}_{j}$ are determined by the lens equation, which for the PML model is given by $\mathbf{y}=\mathbf{x}-\mathbf{x}/|\mathbf{x}|^{2}$ [84]. For $\mathbf{y}=(y,0)$, with $y>0$, the two images on the lens plane $\mathbf{x}_{1,2}=(x_{1,2},0)$ are determined by $x_{1,2}=(y\pm\sqrt{y^{2}+4})/2$, which leads to the following formula for the transmission factor [42]:
$F_{GO}=\left(\sqrt{\mu_{+}}+\sqrt{\mu_{-}}\,e^{2\mathrm{i}\alpha}\right)e^{\mathrm{i}\phi_{1}},$ (10)
Here, the magnification for each image, expressed as $\mu_{\pm}=(v+v^{-1}\pm 2)/4$ with $v\equiv y/\sqrt{y^{2}+4}$, is ultimately a function of $y$ alone, while the phase $\alpha=\pi ft_{M}\tau_{21}-\pi/4$, where $\tau_{21}=2v/(1-v^{2})+\ln[(1+v)/(1-v)]$, also depends on frequency and lens mass. The function $t_{M}\tau_{21}$ represents the time delay between the two images, while $\phi_{1}$ denotes the phase of the first image, taken as a reference [61].
For practical purposes and to minimize excessive computational time, we employ a hybrid function for the transmission factor. This function utilizes the full wave $F$ from Eq. (6) at low frequencies and its GO limit (10) at high frequencies. The simplicity of the GO formula significantly accelerates computations. To ensure a smooth transition between these two solutions, we choose the frequency of the third oscillation maximum, $f_{m}=2.25/(t_{M}\tau_{21})$, as the matching point (see Refs. [61, 65] for details).
III Conditional Variational autoencoder
Given a noisy GW waveform, $h$, with some parameters, $\Lambda$, the objective of Bayesian parameter estimation is to find the posterior probability distribution of the parameters conditioned by this $h$, $p(\Lambda|h)$. Following Bayes theorem, this is given by
$p(\Lambda|h)\propto p(h|\Lambda)p(\Lambda),$ (11)
where $p(h|\Lambda)$ represents the likelihood and $p(\Lambda)$ the prior of the parameters. The common way of approaching this problem with Bayesian inference methods involves several explicit likelihood evaluations, which results in large computational times [95]. In this context, likelihood-free, machine learning methods have been proposed as a way of speeding up the process [75, 76]. A kind of models that have proved specially useful for the estimation of posterior probability distributions are CVAEs. Autoencoders [96] are a type of neural network used to learn efficient, compressed representations of data in a space – the so-called latent space – of lower dimension than the input. An autoencoder consists of two main parts:
  1. 1.
    Encoder: Maps the input data into a lower-dimensional space.
  2. 2.
    Decoder: Reconstructs the original data from the compressed representation.
Figure 1: Configuration of the CVAE network for the training phase. More information about the training flow can be found in Appendix A.
Figure 2: Configuration of the CVAE network for the testing phase. More information about the testing flow can be found in Appendix A.
The goal is to minimize the difference between the input and the output, thereby learning meaningful features of the data without supervision. Autoencoders are widely used in tasks like data compression, denoising, and anomaly detection [97]. A special type of autoencoders are Variational Autoencoders (VAEs) [98]. These learn probabilistic latent representations of data. This means that, unlike traditional autoencoders (which map the input to a fixed point in the latent space), VAEs model the latent space as a probability distribution, typically Gaussian. In other words, the features learnt are the distribution parameters – the mean and the standard deviation in the case of a Gaussian. Then, the decoder reconstructs the data by sampling from this latent space. VAEs are often used in generative modeling, anomaly detection, and unsupervised learning [97, 99]. Finally, a CVAE [77] is an extension of the VAE that introduces conditioning information into the model. This conditioning could be labels, categories, or any side information related to the data. In CVAEs, the conditioning information is used both by the encoder and the decoder, for example, by concatenating it with the input data and with the sample taken from the latent space. In summary, the main difference from VAEs is that CVAEs generate outputs that are conditioned on specific information, making them useful for tasks like controlled data generation and structured prediction (e.g. generating images of a specific class).
In this work, we construct a CVAE following the lines drawn in [75], specializing it for the estimation of the two parameters determining the lensing effect of a PML, $\Lambda_{L}=\{t_{M},y\}$. This CVAE consists on a pair encoder-decoder, $r_{\theta_{1}}$ and $r_{\theta_{2}}$, together with a recognition encoder, $q_{\phi}(z|\Lambda,h)$, where $\theta_{1}$, $\theta_{2}$ and $\phi$ represent the set of trainable parameters of each network and $\Lambda=\{d_{L},m_{1},m_{2},t_{M},y\}$ are the parameters fed to the recognition encoder. The small changes done to the network with respect to the one in [75] are detailed in Appendix A, as well as the details concerning the training process. A scheme of the network used for the training phase can be seen in Fig. 1. Correspondingly, a scheme of the test phase is shown in Fig. 2. Notice that the part of conditioning information mentioned in the last paragraph, is played in this model by the GW strain in the time domain, $h(t)$.
IV Data
Figure 3: Example of a whitened, lensed BBH waveform from the datasets. The waveform is shown in red and the noise in blue. In this case, lensing is strong enough to generate two images.
In this study we rely on utilities from the bilby software package [83] in order to generate both lensed and unlensed BBH merger waveforms. The process of generation is as follows: first, we generate noise data for the Hanford (H1) and Livingston (L1) LIGO detectors at a sampling rate of 2048 Hz, following the methodology described in [100]. Specifically, we use Dataset 3, which consists of purely Gaussian noise, sampled according to power spectral densities calculated from real data from the O3a observation run. Then, we sample the physical parameters of a given waveform from the default bilby BBH prior values save for the component masses, which are uniformly sampled in the $[35,80]\leavevmode\nobreak\ M_{\odot}$ interval and forced to satisfy $m_{1}>m_{2}$. We additionally sample the lensing parameters $y\in\left[0.01,3.0\right]$ and $t_{M}\in[0.001,0.2]$ seconds, as well as a target signal-to-noise ratio (SNR) in the interval $\textrm{SNR}\in[8,40]$. The priors for the rest of parameters are the default priors used by bilby. The physical parameters are then passed to the waveform generator, which uses the IMRPhenomXPHM approximant [101]. This waveform generator computes the transmision factor following the approach explained in Section II. Lensing is applied to the resulting frequency-domain waveform according to
$\tilde{h}_{L}(f;\Lambda)=F(f;\Lambda_{L})\cdot\tilde{h}(f;\Lambda_{U})\,,$ (12)
which is a parametrized version of Eq. (2). In this expression $\Lambda_{U}$ denotes the set of source parameters defining the unlensed waveform, while $\Lambda_{L}=\{t_{M},y\}$ represents the parameters governing lensing effects. Moreover, $\Lambda$ denotes the union of these two sets. This waveform is then injected at the center of a randomly chosen time in an 8-second segment of noise data and the luminosity distance is iteratively adjusted so the SNR of the injected signal matches the target. After this, the data is whitened by inverse spectrum truncation. Since whitening causes artifacts near the edges of the data, we keep only the central 4 seconds for the analysis. The resulting time-series and its physical parameters are then saved to an HDF5 file.
The sizes of the training, validation and testing datasets are, respectively, $10^{6}$, $2\times 10^{4}$ and $250$. These datasets do not intersect to ensure that we can monitorize the generalization ability of the model. An example of a whitened, lensed waveform from one of the datasets is shown in Fig. 3.
V Results
We train our CVAE using the architecture detailed in Appendix A, achieving a smooth evolution of the loss function, as shown in Fig. 4 (specific definitions are provided in the appendix). During the initial 100 epochs, while the annealing process is active, the loss shows greater fluctuations. After this period, when annealing ends and the learning rate is set to $10^{-5}$, the loss evolution becomes noticeably smoother. The loss converges to approximately -2.5, with no signs of overfitting, as both training and validation losses remain close throughout.
Figure 4: Train and validation loss evolution during training.
At the end of training, we save the model from the epoch with the best performance, specifically the epoch with the lowest validation loss, and use it for testing. Using the trained CVAE, we generate 8,000 samples for each of the 250 test waveforms, produced following the guidelines defined in Section IV. These samples define the posterior distributions for each parameter, $y$ and $t_{M}$. A probability-probability (p-p) plot comparing the distribution of parameters of the test set with the distribution of the generated posterior samples is shown in Figure 5. A p-p plot compares the cumulative distribution of predicted probabilities to the cumulative distribution of observed data. In this plot, the $x$-axis represents the quantiles of the cumulative probability predicted by the model (e.g. the quantile range 0% to 100%), whereas the $y$-axis represents the observed frequency of true values falling within these predicted quantile ranges. When the plot lines are close to the diagonal line, it indicates that the model’s predicted probability intervals match the observed data well. In other words, for an $X\%$ quantile on the $x$-axis, approximately $X\%$ of the true values fall within that quantile range in the observed data. A p-p plot close to the diagonal line suggests that the model is well-calibrated: the predicted uncertainty aligns well with the actual uncertainty in the data. As Fig. 5 shows, our trained CVAE yields satisfactory results for both model parameters.
Figure 5: Probability-probability (p-p) plot comparing the posterior probabilities predicted by the CVAE with the true values for the two model parameters. Close alignment with the dashed diagonal line indicates good calibration, meaning that the inferred values fall within the predicted posterior probability intervals as expected.
The top panel of Fig. 6 shows the percentage of test waveforms that fall within the 95% confidence intervals of the posterior distributions, arranged by SNR. Notably, these percentages remain consistent across different SNR levels. Approximately 95% of the true values lie within the 95% confidence intervals for both $t_{M}$ and $y$, aligning with the findings in the p-p plot. We also show the average standard deviation of the posterior distributions in the bottom panel of Fig. 6, which is higher for waveforms with lower SNR. Nevertheless, even in the cases of low SNR, the average standard deviations of the posteriors define confidence intervals significantly narrower than the original prior ranges of $y$ and $t_{M}$.
Figure 6: Top panel: Bar plots showing the percentage of true values of $t_{M}$ (left) and $y$ (right) inside the 95% confidence intervals, arranged by SNR. Bottom panel: same as top panel but showing average standard deviation of the posteriors for $t_{M}$ (left) and $y$ (right).
The scatter plots in Figure 7 show the test set waveforms, labeled by their values of $y$ and $t_{M}$. The color code in the dots indicates whether the waveforms are correctly (blue) or incorrectly (red) processed by the CVAE. Notably, the CVAE successfully estimates $y$ and $t_{M}$ even when $t_{M}$ is close to 0 (i.e. when the lensing effect is nearly absent), demonstrating its ability to handle unlensed waveforms. Additionally, the model accurately detects lensing for values of $y>1$, where the effect is weaker. These cases are challenging to identify, and previous work, such as [78], did not even account for waveforms with those $y$ values.
Figure 7: Scatter plots showing the data points (test set waveforms) based on whether their true values of $y$ (top) or $t_{M}$ (bottom) fall within (blue) or outside (red) the 95% confidence intervals defined by the posteriors.
A noticeable accumulation of errors at higher values of $y$ and $t_{M}$ can be observed in the top and bottom panels of Fig. 7, for $y$ and $t_{M}$, respectively. The underlying cause of this trend is uncertain and might be a stochastic effect. It is important to note that these plots indicate whether values fall within or outside the 95% confidence interval, so we expect about 5% of the waveforms to be incorrectly processed. A deeper investigation focused around the specific range of those high values of $y$ and $t_{M}$ may be worth pursuing to clarify the source of the observed trend and will be reported elsewhere.
Finally, we analyze the efficiency of the CVAE to perform the inference of the lens parameters, comparing our timing estimates to those of bilby. In order to do so we perform parameter estimation on 17 different BBH waveforms and measure the inference (wall-clock) time [^2] The computations were carried out at the ICCUB Nyx’s cluster, equipped with dual AMD EPYC 7763 processors, each with 64 cores.. The results are presented in Table 1. The values reported in this table are obtained from those of Table 3 in Appendix A which shows all timing estimates for the entire set of waveforms. We observe that the CVAE is significantly more efficient than Bilby, achieving speeds about 4 orders of magnitude faster (occasionally even 5 orders of magnitude; see Table 3). To accelerate Bilby’s convergence, we use the true values as priors for all parameters except $y$ and $t_{M}$. For $y$ and $t_{M}$, we set uniform priors ranging from 0 to 3 and from 0 to 0.2 seconds, respectively. When the CVAE is used to provide priors for Bilby, the priors for these two lensing parameters are uniform distributions defined by the 95% confidence intervals of the CVAE’s posterior estimates, i.e. $\mu\pm 2\sigma$, where $\mu$ and $\sigma$ represent the posterior mean and standard deviation respectivelt. In this setup, the average runtime is reduced by 47.9% compared to the previous configuration.
CVAE Bilby Bilby+CVAE Saving
(s) (s) (s) (%)
Average 4 54045 28159 47.9
Maximum 4 145063 81271 79.7
Minimum 4 27197 15932 9.7
Table 1: Comparison of the efficiency between our CVAE and Bilby for parameter estimation. The first two columns report the wall-clock time (in seconds) needed to infer the parameters of a single waveform, showing the longest, shortest, and average durations across 17 runs, for both the CVAE and Bilby in standalone mode. The third column shows the runtime of Bilby when using priors provided by the CVAE. The fourth column presents the percentage of time saved in the three cases.
In addtion to significantly speed up the inference, the use of the CVAE, either combining it with bilby or otherwise, does not degrade the accuracy of the parameter estimation obtained by bilby when used in standalone mode. A proof of this assertion is shown in Figure 8 which depicts a corner plot of the posterior distributions of the two lens parameters, for a particular waveform of our sample, obtained with both bilby and the CVAE in standalone mode, plus the corresponding estimation made with the two approaches combined.
Figure 8: Example of parameter estimation with bilby in standalone mode (blue), the CVAE (red) and a combination of both (green) for the same waveform. Notice that the blue and green posteriors practically overlap, showing that providing priors with CVAE does not degrade the accuracy of the estimations made by bilby.
VI Conclusions and outlook
We have studied the suitability of Conditional Variational Autoencoders (CVAE) to conduct parameter inference of microlensed gravitational waves from binary black hole mergers. Specifically, we have trained a CVAE to estimate the lensing parameters of a point mass lens model, namely $y$ (the relative source position) and $t_{M}$ (the characteristic lensing time). Our study has shown that CVAE are quite capable to perform this task, yielding high performance in terms of both accuracy and computational efficiency.
The CVAE model has displayed substantial advantages over traditional Bayesian methods, including a significant reduction in computation time, up to five orders of magnitude faster than bilby. This efficiency makes CVAE an attractive option for real-time or low-latency applications in GW data analysis, addressing one of the main challenges in current microlensing studies, computational cost. Furthermore, our analysis has revealed that incorporating CVAE-generated priors in the inference can further enhance Bayesian frameworks like bilby. This suggests a promising hybrid approach that combines the rapidness of machine learning and the reliability of traditional statistical inference.
Future research can focus on refining the architecture of the CVAE and making it more suitable for real-world applications. First and foremost, it will be essential to simulate the data using detector noise. In addition, the inclusion of normalizing flows, as proposed by [76], may enhance the model’s performance. Furthermore, incorporating spectrograms as extra input data could enable the CVAE to extract additional features, potentially leading to more precise parameter estimates. Finally, trying the model on real data will be an interesting exercise that could potentially bring new insights to the field of gravitational lensing.
Acknowledgements
The authors are grateful to Mick Wright and Sreekanth Harikumar for their valuable support and guidance with the Gravelamps package. We also thank Narenraju Nagarajan and Melissa Lopez Portilla for their helpful comments on the manuscript. This work was supported by the Spanish Ministry of Science and Innovation (grants PID2021-125485NB-C21 and PID2021-125485NB-C22 funded by MCIN/AEI/10.13039/501100011033 and ERDF, ‘A way of making Europe’), along with the Agència de Gestió d’Ajuts Universitaris i de Recerca (AGAUR), Generalitat de Catalunya (grant SGR-2021-01069), and the Generalitat Valenciana (grant CIPROM/2022/49). RBN acknowledges financial support from NTT DATA Spain and expresses gratitude to Prof. Javier Béjar for his guidance on machine learning matters. OGF is supported by an FCT doctoral scholarship (reference UI/BD/154358/2022). Computations have been performed at the Nyx cluster of the ICCUB. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation.
Appendix A Features of the CVAE model
We include in this appendix details on the architecture of our CVAE and on the training and testing processes. Our model follows the general lines drawn in [75] with some minor modifications that we discuss below. The architecture is specified in Table 2. Notice that the output shape in the case of the convolutional 1D layers is given by the formula
$L_{\rm out}=\bigg{\lfloor}\frac{L_{\rm in}-\text{Kernel Size}}{\text{Stride}}+1\bigg{\rfloor},$ (13)
since we use no padding or dilatation. All the layers, except for the last linear layer of the encoder, decoder and recognition networks, have Leaky ReLU as activation function. The weights of the convolutional block are initialized using the He method at the beginning of the training process. With respect to the CVAE presented in [75], minor changes were carried out, most of them concerning the dimensions of the layers. This was motivated by the fact that our input signal is longer and we are estimating only two parameters. We also added an extra convolutional layer in the shared convolutional block. These decisions were taken in part by trial and error, and the final configuration is the one that achieved the best results.
Component Layer Input Shape Output Shape Parameters
Shared Conv Convolutional [4096, 2] [4033, 96] Kernel Size = 64, Stride = 1
Convolutional [4033, 96] [3970, 96] Kernel Size = 64, Stride = 1
Convolutional [3970, 96] [652, 96] Kernel Size = 64, Stride = 6
Convolutional [652, 96] [589, 96] Kernel Size = 64, Stride = 1
Convolutional [589, 96] [140, 96] Kernel Size = 32, Stride = 4
Convolutional [140, 96] [109, 96] Kernel Size = 32, Stride = 1
Convolutional [109, 96] [39, 96] Kernel Size = 32, Stride = 2, Flatten output $\rightarrow[3744,1]$
Encoder Linear [3744, 1] [4096, 1] -
Linear [4096, 1] [2048, 1] -
Linear [2048, 1] [1024, 1] -
Linear [1024, 1] [128, 1] Reshape output $\rightarrow[2,32,2]$ [^3] A multivariate Gaussian with 32 modes. Notice that the dimension of the latent space is 2.
Recognition Linear [3749, 1]  [^4] The output of the convolutional block concatenated with the vector of real parameters of the gravitational wave. [4096, 1] -
Linear [4096, 1] [2048, 1] -
Linear [2048, 1] [1024, 1] -
Linear [1024, 1] [4, 1] Reshape output $\rightarrow[2,2]$ [^5] A multivariate Gaussian in the latent space.
Decoder Linear [3746, 1]  [^6] The output of the convolutional block concatenated with a sample from the latent space defined by the recognition (training) or the encoder (testing) networks. [4096, 1] -
Linear [4096, 1] [2048, 1] -
Linear [2048, 1] [1024, 1] -
Linear [1024, 1] [4, 1] Reshape output $\rightarrow[2,2]$ [^7] A truncated, multivariate Gaussian as final output for parameter estimation.
Table 2: Layer description of the Variational Autoencoder (VAE). The different parts of the CVAE (shared convolutional block, encoder, recognition network and decoder) are disposed as shown in Figure 1.
The training process starts when the time series is passed through the shared convolutional block. The result is, on the one hand, passed through the enconder network, obtaining a multivariate Gaussian distribution with 32 modes. We call it $\mu_{r_{1}}$ following the notation of [75]. On the other hand, the result is concatenated with five of the source and lens parameters $\Lambda=\{d_{L},m_{1},m_{2},t_{M},y\}$ and passed through the recognition network, obtaining another multivariate Gaussian distribution. This is called $\mu_{q}$. From the latter we sample a vector $z$, which is concatenated with the output of the shared convolutional block and passed through the decoder network, obtaining a third multivariate Gaussian distribution, $\mu_{r_{2}}$. With these four elements, $\mu_{q}$, $\mu_{r_{1}}$, $\mu_{r_{2}}$ and $z$, we compute the loss function, which is the sum of two terms:
  1. 1.
    The reconstruction term $(r_{2})$: Computed as the negative log likelihood of the lens parameters $\Lambda_{L}=\{t_{m},y\}$ being a sample of the distribution described by the parameters $\mu_{r_{2}}$. This term is minimized when there is a high probability of sampling $\Lambda_{L}$ from $\mu_{r_{2}}$:
    ${\rm RL}\,[\mu_{2};\Lambda_{L}]=-\log(r_{2}(\Lambda_{L}))$ (14)
  2. 2.
    The KL-divergence: Computed between the distributions defined by the parameters $\mu_{q}$ and $\mu_{r_{1}}$, $q(z|h,\ \Lambda)$ and $r_{1}(z|\Lambda)$. We can approximate the KL-divergence as
    ${\rm KL}\,[q(z|h,\ \Lambda_{L})||r_{1}(z|h)]\approx\log\bigg{(}\frac{q(z|h,\ \Lambda)}{r_{1}(z|h)}\bigg{)}\,.$ (15)
    Since this is equal to $\log(q(z|h,\ \Lambda))-\log(r_{1}(z|h))$, we just have to compute the log probabilities of $z$ being a sample of both distributions. This term attains its minimum value when the distributions $q$ and $r_{1}$ look alike.
All the trainable parameters are then updated during back-propagation, driven by this loss function. We apply the same annealing process for the KL divergence and the truncated Gaussian distribution from the output than in [75]. Both the contribution of said loss term and the range of the truncated Gaussian, are controlled by a parameter $\beta$ that multiplies the KL diverge in the loss function. This parameter increases logarithmically from 0 to 1 between epochs 10 and 20. The lower and upper bounds of the truncated Gaussian vary, respectively, from -10 to 0 and from 11 to 1. In each epoch, the CVAE visits $2\times 10^{4}$ instances, and we perform 1100 epochs. Therefore, the CVAE visits the whole dataset (made out of $10^{6}$ waveform signals) 22 times. The batch size is set equal to 1500 and the learning rate is set equal to $10^{-4}$ during the first 100 epochs and equal to $10^{-5}$ during the rest of the training process. We use an Adam optimizer with L2 regularization (a weight decay of $10^{-5}$) and a validation set of $2\times 10^{4}$ waveforms.
Once the CVAE is trained we can test it on a test set. We use a set of 250 waveforms that have never been seen by the model before to test its generalization capability. During the testing process, the recognition encoder is not used. Instead, the time series is passed through the convolutional block and then through the encoder. A sample is taken from the latent space, concatenated with the output of the convolutional block and passed through the decoder, obtaining the parameters of two truncated Gaussian distributions. By passing the same time series $N$ times through the model, one gets $N$ different pairs of truncated gaussians. We take a sample from each of these distributions, obtaining a set of $N$ samples that define a posterior probability that can be represented in a corner plot. When $N=8000$ this process takes, on average, around 4 seconds per waveform. The GPU used for training and testing is an NVIDIA A40.
As a final test, we compare the efficiency of the CVAE with bilby [83] for parameter estimation. We also study if using the CVAE to provide priors to bilby helps reducing the computational time of the inference. The results of this analysis is shown in Table 3.
CVAE Bilby Bilby + CVAE Time saved SNR $y$ $t_{M}$
(s) (s) (s) % (s)
4 62374 30370 51.31 18.14 0.88 0.046
4 36327 29156 19.74 25.84 2.97 0.072
4 102661 36330 64.61 24.49 0.74 0.184
4 47179 22119 53.12 28.77 1.33 0.033
4 27197 21376 21.40 16.88 2.48 0.129
4 47195 18809 60.15 28.00 0.45 0.147
4 36930 15932 56.86 38.42 1.40 0.131
4 145063 81271 43.98 15.28 1.86 0.002
4 48646 29838 38.66 18.62 1.12 0.088
4 82371 35772 56.57 8.98 0.65 0.097
4 34325 25982 24.31 27.66 2.16 0.092
4 74097 33421 54.90 22.26 0.34 0.086
4 45966 29340 36.17 13.26 0.44 0.150
4 69607 32147 53.82 24.31 1.10 0.042
4 134742 27308 79.73 32.64 0.61 0.081
4 29511 26648 9.70 10.74 2.84 0.119
4 29263 18671 36.20 31.37 1.48 0.077
Table 3: Comparison of inference times (in seconds) for the CVAE and bilby. Each row represents a different waveform, with the corresponding SNR and $y$ and $t_{M}$ values